OpenStudy (anonymous):

,@melissaneedshelp is a math tutor in the academic center. She charges $210 for the first 7 hours of tutoring and $20 for each additional hour of tutoring. If she earned $410 tutoring 1 student, how many total hours did she spent tutoring the student?

OpenStudy (anonymous):

17 hours

Parth (parthkohli):

210 + 20x = 410 => 20x = 410 - 210 => 20x = 200 => x = 200/20 => x = 10 So 7 hours plus 10 additional hours = 17 hours :)

OpenStudy (anonymous):

gud question!

OpenStudy (anonymous):

It is indeed!